# Quokka-based understanding of outflows (QED) - II. X-ray metallicity gradients as a signature of galactic wind metal loading

### v1

@@ Abstract
pubyear: 2024pagerange: Quokka-based understanding of outflows (QED) - II. X-ray metallicity gradients as a signature of galactic wind metal loadingQuokka-based understanding of outflows (QED) - II. X-ray metallicity gradients as a signature of galactic wind metal loading
@@ 1 Introduction
Metals are synthesized in the cores of stars that, at least in spiral galaxies such as the Milky Way, reside in a thin disc. However, observations reveal that these metals are not confined to the disc. Instead, $\sim 50\%$ of the mass of metals like oxygen is found in galactic halos, at distances as large as $\sim 100\ \text{kpc}$ from the central galaxy (Peeples et al., 2014). This raises the intriguing question of how metals are transported from the disc to the halo. The most plausible explanation is that they are carried by galactic outflows, powerful winds driven by supernova (SN) feedback that can transport material, momentum and energy from the disc to the halo.
@@
Physically, these outflows are driven by SN thermal energy injection that heats up the surrounding gas, driving it to expand and break out of the disc. Once hot bubbles break out, the momentum-carrying gas expands freely and eventually ends up in the circum-galactic medium (CGM) around the galaxy. Because this hot gas is generated as a direct result of SN activity, it is expected to be loaded with metals and therefore enriches the CGM with metals. However, the balance in the gas escaping the galaxy between unmixed supernova ejecta and swept-up ambient gas from the galactic interstellar medium (ISM) – which is also metal-enriched but much less so than the fresh supernova ejecta – is highly uncertain.
One potential way to quantify the fraction of SN-injected metals carried by the outflowing gas is by studying the diffuse X-ray emission around star-forming galaxies. Such emission is linked to star-formation processes occurring within the disc (Strickland et al., 2004; Grimes et al., 2005; Tüllmann et al., 2006; Owen & Warwick, 2009; Li & Wang, 2013; Wang et al., 2016), and is useful for studying metallicity due to the large number of abundance-diagnostic lines that appear in soft X-rays. In particular, the starburst M82 serves as a textbook case of SN-driven outflows and much attention has been given to measuring the metallicity of the outflows launched from the central region in the X-ray (e.g., Read & Stevens, 2002; Origlia et al., 2004; Ranalli et al., 2008; Konami et al., 2011; Lopez et al., 2020; Fukushima et al., 2024), with complementary efforts at other wavelengths (e.g., Cameron et al., 2021).
@@
All these studies have revealed that the outflows M82 have non-uniform metallicities. Read & Stevens (2002) and Origlia et al. (2004) found that the outflows are enhanced in $\alpha/\mathrm{Fe}$ relative to the disc, while Lopez et al. (2020) measured the gradients of metallicity along the minor axis of M82 and reported a clear trend of declining abundances and gas temperatures with distance away from the central galaxy for several $\alpha$ elements. They hypothesize that the outflow profile of M82 is a result of mixing between different temperature phases and not simply because of the adiabatic expansion of hot wind. However, the nature of X-ray emission poses challenges for interpretation. For example, spectral fitting relies on simplifying assumptions about temperature profiles of X-ray emitting gas (Vijayan & Li, 2022). One way to overcome such limitations is to provide a theoretical basis for the interpretation of metallicity gradients using simulations.
Numerous simulations have studied the properties of metals in galactic outflows (e.g., Melioli et al., 2013; Emerick et al., 2018; Andersson et al., 2023; Rey et al., 2024; Kim et al., 2020; Schneider et al., 2020; Steinwandel et al., 2024). However, these studies have been limited in their scope as they focus on theoretical estimates of metal loading and its gradients. While metal loading of the outflows is closely related to how the metallicity gradients are established, it is important to reproduce the trends using steps that are near-identical to those used in observations as fitting procedures might alter the theoretically expected trends.
@@
Moreover, correctly capturing the properties of outflows, and particularly the distributions of metals within them, is an extraordinarily challenging numerical problem due to its simultaneous requirements of very high resolution and very large volumes. Resolution is required because mixing between the hot gas that emits in X-rays and the cooler gas that can be studied in optical or radio is a critical process in determining both the temperature and the metal content of the hot phase. The rate of mixing between the hot and cooler phases is easily affected by poor resolution, with Lagrangian methods tending to under-mix without an explicit sub-grid diffusion model (Shen et al., 2010; Escala et al., 2018), and Eulerian methods tending to over-mix due to numerical diffusion. At a minimum, no simulation that has to rely on subgrid models of SN explosions to avoid over-cooling can reliably study metal mixing, a condition that rules out even zoom-in cosmological simulations, and even properly resolving the Sedov-Taylor phase may be insufficient to yield converged results for clustered SNe (Gentry et al., 2019). At the same time, large volumes are needed because in order to replicate trends in observables we require large volumes that ensure that the winds have developed fully and that we can trace them out to the multi-kpc distances away from the disc that are typical of observational studies. Few existing simulations meet these simultaneous requirements.
@@
In this work, we use QED [^1] QUOKKA-based Understanding of Outflows Derived from Extensive, Repeated, Accurate, Thorough, Demanding, Expensive, Memory-consuming, Ongoing Numerical Simulations of Transport, Removal, Accretion, Nucleosynthesis, Deposition, and Uplifting of Metals (QUOD ERAT DEMONSTRANDUM, or QED) simulations, introduced in Vijayan et al. (2024, hereafter QED I), to study metals in galactic winds with mock X-ray observations. QED is a suite of 3D hydrodynamic simulations of galactic winds generated by the new code Quokka (Wibking & Krumholz, 2022; He et al., 2024), a state-of-the-art adaptive mesh refinement (AMR) radiation-hydrodynamics (RHD) tool optimized for graphics processing units (GPUs). The simulation we use follows SN-driven outflows with a uniformly high resolution of $2$ pc that allows us to accurately follow the metal exchanges taking place in the multi-phase outflows, and that QED I demonstrate via convergence studies is sufficient to capture their properties. At the same time, the simulation domain extends to ±$\pm 4$ kpc above and below the galactic midplane, enabling us to follow trends up to $\sim$ a few kpc above the disc.
@@
This paper is organized as follows. Section 2 provides a brief summary of the setup of the QED simulations, and the method we use to generate synthetic observations from them and then analyse them using the same methods that are routinely applied to real observations. Section 3 presents our main results related to fits of the mock observations, and Section 4 discusses the interpretation of the results and their implications. We outline our major conclusions in Section 5.
@@ 2.1 The QED simulations
QED is a suite of fixed-grid simulations of a $\sim 1\times 1$ kpc2 patch of the Milky Way disc; QED I presents results and a convergence study for Solar neighbourhood conditions, while Vijayan et al. (2025) extends the model grid to a wide range of galactic environments. In this paper, we use the highest-resolution case presented in QED I, and we refer readers to that paper for full details of the numerical setup, which we only summarise here. The simulation has a simulation box centered on the disc midplane that extends to ±$\pm 4$ kpc and a uniform resolution of $512\times 512\times 4096$, giving a grid spacing of $2$ pc. The gas is confined by the gravity of the stellar disc and dark matter halo and driven by SNe, which occur at a constant rate calibrated to the Kennicutt (1998) relation and distributed exponentially in height around the disc with a scale height of 150 pc. SNe are treated as injections of $10^{51}$ erg of pure thermal energy; the resolution is high enough that the Sedov-Taylor phase is well-resolved, and no subgrid models are required. Gas cooling is implemented through a tabulated cooling function derived from the Grackle cooling model (Smith et al., 2017). After nearly $100$ Myr of SN feedback, the simulation reaches quasi-steady-state outflow rates. We show $y=0$ slices of density and temperature from a snapshot at $110$ Myr in the top two panels of Figure 1. The outflows clearly possess a multi-phase structure characterised by a broad range of temperatures from $<10$ K to $>10^{7}$ K, with high (low) temperature gas being associated with low (high) densities.
@@
Metals in the simulations can be either injected by SNe or present in the initial conditions. For the former, each SN injects a fixed mass $\Delta M_{\mathrm{SN}}=1$ M of oxygen into the same cell in which it deposits thermal energy, which is subsequently advected as a passive scalar. The second source of metals is the initial background metallicity of the ISM present at the start of the simulation, which we parameterise by $Z_{\rm bg}$. We scale it to Solar abundances, so the initial oxygen mass density in a cell of total mass density $\rho$ is $\rho_{\mathrm{O}}=\rho(Z_{\mathrm{bg}}/Z_{\odot})Z_{\mathrm{O,\odot}}$, where $Z_{\odot}$ denotes Solar metallicity and we set the Solar oxygen abundance to $Z_{\mathrm{O,\odot}}=8.6\times 10^{-3}$. We treat $Z_{\rm bg}$ as an independent parameter that we can alter in the post-processing of the simulations. [^2] We note here that in QED I all gas cools at Solar metallicity, irrespective of the local metallicity. This is the downside to the flexibility we gain by our purely passive treatment of metals. The total metal density of a cell is the sum of the background and SN-added contributions – see Section 2.3.1 of QED I for full details of the computation. Once we have computed the oxygen metallicity in a cell, we assume Solar-scaled abundances for all other elements. This assumption is not entirely realistic, since we are modeling only type II SNe that produce predominantly $\alpha$ elements rather than type Ia SN that produce iron peak elements, and thus our added metal field should have a different $\alpha$/Fe ratio than our background field; we discuss the implications of this further below.
@@ 2.2.1 Generating photons using pyXSIM
The pyXSIM package generates a set of X-ray photons from each cell in the simulation under the assumption of a collisionally ionized and optically thin plasma, following the Astrophysical Plasma Emission Code (APEC; Smith et al., 2001). Because we are interested in the diffuse X-ray emission from galactic winds, we limit the photon energies to the soft X-ray band, with energies $0.4-2.0$ keV; this is the energy range over which Chandra is the most sensitive, contains most (but not all) of the important diagnostic lines for metallicity, and in practice is usually the only usable band for spatial studies of galactic winds because emission from the relatively harder X-ray band ($2-5$ keV) is significant compared to the soft X-ray contribution only in regions close to the disc, and drops sharply away from it (Vijayan et al., 2018).
Apart from gas density and temperature, pyXSIM also requires the metallicity. As discussed in Section 2.1, this depends on the choice of initial background metallicity $Z_{\mathrm{bg}}$; we carry out our analysis for the cases $Z_{\mathrm{bg}}=0$, 0.2, 0.5, 1, and 2, providing a wide baseline. We summarise our choices for the other parameters required by pyXSIM, including the exposure time, collecting area, and distance to source, in Table 1. Our choice of collecting area is typical of Chandra, and our distance is roughly the distance to M82. In order to get a “pure” unadulterated spectrum, we disable absorption, backgrounds, and foregrounds.
One parameter choice that requires some comment is the exposure time, for which we choose for now an unrealistically-large value of $1000$ Ms. We require such a large exposure time due to the limits of our simulation, which includes only a $1\,\mathrm{kpc}\times 1\,\mathrm{kpc}$ region that represents a small fraction of an actual Milky Way-sized galaxy (typical size closer to $\sim 10\,\mathrm{kpc}\times 10\,\mathrm{kpc}$). [^3] Conversely, our simulation volume would be more realistic for a circumnuclear starburst such as M82, but in such a galaxy the wind would be much denser and more powerful than in our Solar neighbourhood setup, and thus much brighter in the X-rays. In order to produce high signal-to-noise spectra we need to generate a proportionately larger number of photons, which is why we choose such a large exposure time. We will consider the effects of smaller and more realistic signal-to-noise ratios (SNRs) in Section 3.4.

### v2

@@ Abstract
pubyear: 2025pagerange: Quokka-based understanding of outflows (QED) - II. X-ray metallicity gradients as a signature of galactic wind metal loadingQuokka-based understanding of outflows (QED) - II. X-ray metallicity gradients as a signature of galactic wind metal loading
@@ 1 Introduction
Metals are synthesized in the cores of stars that, at least in spiral galaxies such as the Milky Way, reside in a thin disc. However, observations reveal that these metals are not confined to the disc. Instead, $\sim 50\%$ of the mass of metals like oxygen is found in galactic halos, at distances as large as $\sim 100\ \text{kpc}$ from the central galaxy (Tumlinson et al., 2011; Peeples et al., 2014). This raises the intriguing question of how metals are transported from the disc to the halo. The most plausible explanation is that they are carried by galactic outflows, powerful winds driven by supernova (SN) feedback that can transport material, momentum and energy from the disc to the halo.
@@
Physically, these outflows are driven by SN thermal energy injection that heats up the surrounding gas, driving it to expand and break out of the disc. Once hot bubbles break out, the momentum-carrying gas expands freely and eventually ends up in the circum-galactic medium (CGM) around the galaxy. Because this hot gas is generated as a direct result of SN activity, it is expected to be loaded with metals and therefore enriches the CGM with metals. However, the balance in the gas escaping the galaxy between unmixed supernova ejecta and swept-up ambient gas from the galactic interstellar medium (ISM) – which is also metal-enriched but much less so than the fresh supernova ejecta – is highly uncertain. It is also highly important: the fraction of SN-produced metals that are promptly lost to the outflow rather than being retained in the disc is a crucial free parameter in models of galactic chemical evolution (e.g., Peeples & Shankar, 2011; Telford et al., 2019; Forbes et al., 2019; Johnson & Weinberg, 2020; Sharda et al., 2021; Kravtsov & Manwadkar, 2022).
@@
One potential way to quantify the fraction of SN-injected metals carried by the outflowing gas is by studying the diffuse X-ray emission around star-forming galaxies. Such emission is linked to star-formation processes occurring within the disc (Strickland et al., 2004; Grimes et al., 2005; Tüllmann et al., 2006; Owen & Warwick, 2009; Li & Wang, 2013; Wang et al., 2016), and is useful for studying metallicity due to the large number of abundance-diagnostic lines that appear in soft X-rays. In particular, the starburst M82 serves as a textbook case of SN-driven outflows and much attention has been given to measuring the metallicity of the outflows launched from the central region in the X-ray (e.g., Read & Stevens, 2002; Origlia et al., 2004; Ranalli et al., 2008; Konami et al., 2011; Lopez et al., 2020; Fukushima et al., 2024). All these studies have revealed that the outflows M82 have non-uniform metallicities. Read & Stevens (2002) and Origlia et al. (2004) found that the outflows are enhanced in $\alpha/\mathrm{Fe}$ relative to the disc, while Lopez et al. (2020) measured the gradients of metallicity along the minor axis of M82 and reported a clear trend of declining abundances and gas temperatures with distance away from the central galaxy for several $\alpha$ elements. They hypothesize that the outflow profile of M82 is a result of mixing between different temperature phases and not simply because of the adiabatic expansion of hot wind. However, the nature of X-ray emission poses challenges for interpretation. For example, spectral fitting relies on simplifying assumptions about temperature profiles of X-ray emitting gas (Vijayan & Li, 2022). One way to overcome such limitations is to provide a theoretical basis for the interpretation of metallicity gradients using simulations.
While M82 is the best-studied case, thanks to its proximity and nearly edge-on geometry, there is also substantial evidence for metallicity variations in the winds of other galaxies, and in phases other than the hot, X-ray-emitting one. Lopez et al. (2023) find a metallicity gradient in the hot phase of the wind of NGC 253 that is similar to that in M82. Cameron et al. (2021) identify a clear negative metallicity gradient along the minor axis of the wind of edge-on starburst galaxy Markarian 1486 using direct (electron temperature-based) measurements of metallicity in the warm ionised phase. Martin et al. (2002) report significant $\alpha$/Fe enhancement in the hot wind phase of the dwarf starburst NGC 1569, but Hamel-Bravo et al. (2024) find no comparable enhancement in the warm ionised gas, suggesting a composition difference between these two components. Xu et al. (2022) also conclude that there is likely a significant difference between the metal content of the warm ionised and hot components of the wind in their absorption spectroscopic study of 45 low-redshifts starbursts drawn from the CLASSY sample (Berg et al., 2022). Chisholm et al. (2018) carry out absorption spectroscopy in a sample of seven local galaxies and find evidence that metallicities in the warm ionised wind phase are significantly enhanced compared to the interstellar media of the galaxies driving those winds. Significantly, their sample contains ordinary star-forming galaxies in addition to starbursts, and there is no significant difference in the metal enhancement between the two groups. This suggests that, while most studies of metal-enriched winds to date have focused on starbursts simply because their winds are bright and easy to observe, metal-enrichment is a phenomenon common to all galactic winds, not just the winds of starbursts.
There have also been numerous attempts to use simulations to study the metal content of galactic winds. This problem is generally impossible to address in cosmological simulations, which lack the resolution to capture the phase structure of winds without relying on sub-grid models (e.g., Smith et al., 2024). However, even excluding cosmological simulations, there are a range of approaches that represent different trade-offs between volume, resolution, and physical complexity. At the finest resolution but lowest volume end are “tall-box” simulations (e.g., Kim et al., 2020; Rathjen et al., 2021) that simulate only a portion of a galactic disc, while at the larger-volume, lower-resolution end are simulations of entire isolated dwarf galaxies (e.g., Melioli et al., 2013; Emerick et al., 2018, 2019, 2020; Schneider et al., 2018, 2020; Andersson et al., 2023; Rey et al., 2024; Steinwandel et al., 2024b, a). Similarly, there are a range of levels of physical complexity, with some simulations using relatively simple treatments of ISM physics in order to maximise resolution (e.g., Melioli et al., 2013; Schneider et al., 2018, 2020) while others include live treatments of physical processes such as radiative transfer (e.g., Rey et al., 2024) and cosmic rays (e.g., Rathjen et al., 2021) or including multiple elements with different nucleosynthetic origins (e.g., Emerick et al., 2018, 2019, 2020). Moreover, recent high-resolution dwarf galaxy simulations have incorporated multiple feedback channels (e.g., Smith et al., 2021; Gutcke et al., 2021; Deng et al., 2024) or focused on star-by-star physics in isolated systems (e.g., Hu et al., 2016, 2017; Hu, 2019; Hu et al., 2023a, b), highlighting the importance of resolving supernova remnants and properly modeling radiation in shaping multiphase, metal-enriched outflows. However, there have thus far been limited efforts to carry out detailed observational comparisons between these simulations and X-ray observations; the primary exception is the work of Schneider & Mao (2024), who generate detailed synthetic X-ray maps from their simulations. However, those simulations do not include metals, and thus are not directly usable for comparing to observations of metallicity variation in winds such as those summarised above.
@@
Part of the reason for this relative paucity of studies of the metal content of winds including detailed observational comparisons is the numerical difficulty of the problem. Simulations targeting these observations require very high resolution because mixing between the hot gas that emits in X-rays and the cooler gas that can be studied in optical or radio is a critical process in determining both the temperature and the metal content of the hot phase. The rate of mixing between the hot and cooler phases is easily affected by poor resolution, with Lagrangian methods tending to under-mix without an explicit sub-grid diffusion model (Shen et al., 2010; Escala et al., 2018), and Eulerian methods tending to over-mix due to numerical diffusion. At a minimum, no simulation that has to rely on sub-grid models of SN explosions to avoid over-cooling can reliably study metal mixing, a condition that rules out even zoom-in cosmological simulations, and even properly resolving the Sedov-Taylor phase may be insufficient to yield converged results for clustered SNe (Gentry et al., 2019). At the same time, observations of wind metal content generally target regions extending to several kpc off galactic discs, and thus simulations attempting to compare to these observations must have similar extents. Few existing simulations meet these simultaneous requirements.
@@
In this work, we use QED [^1] QUOKKA-based Understanding of Outflows Derived from Extensive, Repeated, Accurate, Thorough, Demanding, Expensive, Memory-consuming, Ongoing Numerical Simulations of Transport, Removal, Accretion, Nucleosynthesis, Deposition, and Uplifting of Metals (QUOD ERAT DEMONSTRANDUM, or QED) simulations, introduced in Vijayan et al. (2024, hereafter QED I), to study metals in galactic winds with mock X-ray observations. QED is a suite of 3D hydrodynamic simulations of galactic winds generated by the new code Quokka (Wibking & Krumholz, 2022; He et al., 2024), a state-of-the-art adaptive mesh refinement (AMR) radiation-hydrodynamics (RHD) tool optimized for graphics processing units (GPUs). In terms of the taxonomy of simulation approaches discussed above, the QED suite lies at the maximum resolution end of the spectrum: tall box rather than whole-galaxy, and including a relatively simple treatment of ISM physics and feedback that includes only driving by randomly-distributed SNe. QED’s primary virtue is the very large dynamic range afforded by GPU acceleration: the simulations feature a uniformly high resolution of $2$ pc that allows us to accurately follow the metal exchanges taking place in the multi-phase outflows. QED I demonstrates that, at the 2 pc resolution achieved in the QED suite, the wind mass flux, metal flux, and metallicity as a function of height off the midplane are all numerically converged. At the same time, the simulation domain extends to ±$\pm 4$ kpc above and below the galactic midplane, enabling us to follow trends up to a few kpc above the disc, capturing the off-disc regions typically targeted in observations. Of course QED – like all tall box simulations – still captures only a small fraction of the volume of a real galactic wind in a galaxy like the Milky Way, which has an area much larger than 1 kpc2 area of the QED tall box. However, it is reasonable to hypothesize that the QED simulation box is statistically representative of the full wind, and is therefore suitable for use in simulated observations provided that we compensate for the smaller volume. Finally, the Quokka code on which the QED simulations are based uses a very high-resolution hydrodynamics scheme based on a PPM solver with a monotonized-central (MC) limiter, using a combination of a dual energy formalism and first-order flux correction to maintain stability in problematic cells (Wibking & Krumholz, 2022); this scheme should be significantly less dissipative, and thus better able to follow phase and metal mixing, than the more dissipative methods used in some earlier simulations.
@@
This paper is organized as follows: Section 2 provides a brief summary of the setup of the QED simulations, and the method we use to generate synthetic observations from them and then analyse them using the same techniques that are routinely applied to real observations. Section 3 presents our main results related to fits of the mock observations, and Section 4 discusses the interpretation of the results and their implications. We outline our major conclusions in Section 5.
@@ 2.1 The QED simulations
QED is a suite of fixed-grid simulations of a $\sim 1\times 1$ kpc2 patch of the Milky Way disc; QED I presents results and a convergence study for Solar neighbourhood conditions, while Vijayan et al. (2025) extends the model grid to a wide range of galactic environments. In this paper, we use the highest-resolution case presented in QED I, and we refer readers to that paper for full details of the numerical setup, which we only summarise here. The simulation has a simulation box centered on the disc midplane that extends to ±$\pm 4$ kpc and a uniform resolution of $512\times 512\times 4096$, giving a grid spacing of $2$ pc, and is run for 116 Myr, by which point the outflow has reached statistical steady state. We show mass outflow rates and loading factors as a function of time and simulation resolution to demonstrate convergence in Appendix A. The gas is confined by the gravity of the stellar disc and dark matter halo and driven by SNe, which occur at a constant rate calibrated to the Kennicutt (1998) relation and distributed exponentially in height around the disc with a scale height of 150 pc. SNe are treated as injections of $10^{51}$ erg of pure thermal energy; the resolution is high enough that the Sedov-Taylor phase is well-resolved, and no sub-grid models are required. Gas cooling is implemented through a tabulated cooling function derived from the Grackle cooling model (Smith et al., 2017). After nearly $100$ Myr of SN feedback, the simulation reaches quasi-steady-state outflow rates. We show $y=0$ density and temperature slices from a snapshot in $110$ Myr in the top two panels of Figure 1. The outflows clearly possess a multi-phase structure characterised by a broad range of temperatures from $<10$ K to $>10^{7}$ K, with high (low) temperature gas being associated with low (high) densities.
@@
Metals in the simulations can be either injected by SNe or present in the initial conditions. For the former, each SN injects a fixed mass $\Delta M_{\mathrm{SN}}=1$ M of oxygen into the same cell in which it deposits thermal energy, which is subsequently advected as a passive scalar. The second source of metals is the initial background metallicity of the ISM present at the start of the simulation, which we parameterise by $Z_{\rm bg}$. We scale it to Solar abundances, so the initial oxygen mass density in a cell of total mass density $\rho$ is $\rho_{\mathrm{O}}=\rho(Z_{\mathrm{bg}}/Z_{\odot})Z_{\mathrm{O,\odot}}$, where $Z_{\odot}$ denotes Solar metallicity and we set the Solar oxygen abundance to $Z_{\mathrm{O,\odot}}=8.6\times 10^{-3}$. We treat $Z_{\rm bg}$ as an independent parameter that we can alter in the post-processing of the simulations. [^2] We note here that in QED I all gas cools at Solar metallicity, irrespective of the local metallicity. This is the downside to the flexibility we gain by our purely passive treatment of metals. The total metal density of a cell is the sum of the background and SN-added contributions – see Section 2.3.1 of QED I for full details of the computation. Once we have computed the oxygen metallicity in a cell, we assume Solar-scaled abundances for all other elements. This assumption is not entirely realistic, since we are modeling only type II SNe that produce predominantly $\alpha$ elements rather than type Ia SN that produce iron peak elements, and thus our added metal field should have a different $\alpha$/Fe ratio than our background field; we discuss the implications of this in more detail below.
@@ 2.2.1 Generating photons using pyXSIM
The pyXSIM package generates a set of X-ray photons from each cell in the simulation under the assumption of a collisionally ionized and optically thin plasma, following the Astrophysical Plasma Emission Code (APEC; Smith et al., 2001). Because we are interested in the diffuse X-ray emission from galactic winds, we limit the photon energies to the soft X-ray band, with energies $0.4-2.0$ keV; this is the energy range over which Chandra is the most sensitive, contains most (but not all) of the important diagnostic lines for metallicity, and in practice is usually the only usable band for spatial studies of galactic winds because emission from the relatively harder X-ray band ($2-5$ keV) is significant compared to the soft X-ray contribution only in regions close to the disc, and drops sharply away from it (Vijayan et al., 2018). One complication in generating synthetic observations from our simulation is that we simulate only a 1 kpc2 patch that represents a small fraction of an actual Milky Way-sized galaxy (typical area closer to 100 kpc2). [^3] Conversely, our simulation volume would be more realistic for a circumnuclear starburst such as M82. However, in such a galaxy the wind would be much denser and more powerful than in our Solar neighbourhood setup, and thus much brighter in the X-rays. In order to compensate for this, we tile 100 copies of our simulation box together in order to generate our synthetic spectra – that is, we generate 100 times as many X-ray photons from each cell per unit time as would be expected, on the basis that a realistically-large galaxy would contain 100 times as many cells with similar physical conditions.
In addition to gas density and temperature, pyXSIM also requires metallicity. As discussed in Section 2.1, this depends on the choice of initial background metallicity $Z_{\mathrm{bg}}$; we carry out our analysis for the cases $Z_{\mathrm{bg}}=0$, 0.2, 0.5, 1, and 2, providing a wide baseline. We summarise our choices for the other parameters required by pyXSIM, including the exposure time, collecting area, and distance to source, in Table 1. Our choice of collecting area is typical of Chandra, and our distance is roughly the distance to M82, and is similar to the distances of other nearby galaxies whose wind metallicities have been measured in X-rays (e.g., NGC 1569 – Martin et al. 2002; NGC 253 – Lopez et al. 2023). In order to get a “pure” unadulterated spectrum, we disable absorption, backgrounds, and foregrounds, and we use a large exposure time of 10 Ms. We will consider the effects of smaller and more realistic signal-to-noise ratios (SNRs) in Section 3.4. We do not seek to explore the effects of confusion due to absorption, backgrounds, and foregrounds because these are all highly variable across the sky, and thus it is not possible to make general statements about the extent to which they can be corrected and the level of observational uncertainty introduced by the need to make such corrections.
@@
2.2.2 Characteristics of the X-ray emitting gas
Before proceeding to simulated observations, it is important to examine the characteristics of the X-ray emitting gas, both to ensure that they are physically reasonable and to ensure that we adequately resolve the emitting regions in our simulations. The reason that resolution is a potential concern is that, thanks to the strong density and temperature dependence of X-ray emissivity, in a multiphase medium such as our simulated wind, X-ray emission can be dominated by relatively thin interface layers between the hot volume-filling phase and the cooler entrained gas that can be hard to resolve (Toalá & Arthur, 2018; Vijayan et al., 2018).
To this end, we examine the cumulative distribution functions (CDFs) of X-ray emission with respect to gas temperature $T$ and volume $V$ for the same snapshot (and using the same $Z_{\mathrm{bg}}=0.5Z_{\odot}$ as shown in Figure 1. For the purpose of analysis discussed here and in subsequent Sections, we partition the simulation domain into 11 regions, mirroring the approach used by Lopez et al. (2020); we mark our regions in the bottom panel of Figure 1. The central region, labeled ‘D’, corresponds to the area where SNe are injected and extends up to a height of 489.03 pc (i.e., 256 cells), while the remaining 10 regions (‘A/B1’ to ‘A/B5’) cover the outflow zone, each spanning 733.55 pc (i.e., 384 cells) in height.
We show the CDFs we obtain for each region in Figure 2. The two panels of this Figure show the fraction of total X-ray emission in the 0.4-2.0 keV band arising from gas with temperature $<T$ (left panel) and from a a given fraction of total volume (right). As expected based on the models of Toalá & Arthur (2018), the left panel of this Figure shows that X-ray emission in our simulation is dominated by moderately ($\sim$ few $10^{6}$ K) rather than extremely ($\gtrsim 10^{7}$ K) hot gas. The right panel, however, shows that the volume occupied by this gas is substantial, and is thus well-resolved within our simulations. Quantitatively, except in region ‘D’, we find that $\approx 50\%$ of the emission emerges from 30% of the volume. Given that our simulation contains a total of $\approx 10^{9}$ computational cells, even 30% of this number constitutes very good resolution: quantitatively, 30% of the volume corresponds to a resolution of roughly $685^{3}$ cells. Resolution is slightly worse in region ‘D’, where 50% of the emission emerges from just under 10% of the volume, but even 10% of our simulation volume corresponds to $475^{3}$ cells. The lower volume filling fraction in region ‘D’ is not surprising since this region is dominated by fresh SNe ejecta with very high metallicities. We therefore conclude that the emitting regions in our simulations are relatively well-resolved.
[…]
Figure 2: Cumulative distribution of X-ray emission in the 0.4–2.0 keV band with respect to gas temperature (left) and volume (right). Different colours correspond to different regions within the simulation, as explained in the main text; we show only the ‘A’ regions at positive $z$ and omit the ‘B’ regions at negative $z$ to reduce clutter in the plot, but the results for the ‘A’ and ‘B’ regions are qualitatively identical. We see that the X-ray emitting gas lies within a narrow temperature range ($\approx 10^{6}-10^{7}$ K), but that the emitting regions are reasonably well-resolved, with $50\%$ of the emission coming from $\approx 30\%$ of the volume in all regions but D, where the emission arises from $\approx 10\%$ of the volume.